# Functionally Constrained Algorithm Solves Convex Simple Bilevel Problems

### v1

@@ Appendix C Proofs for Section 5
$\epsilon_{g}=g({\bf{x}}_{T})-g^{*}\geq\frac{C_{g}D}{2(1+\sqrt{T})},$
which implies that any zero-respecting first-order methods need at least
$T=\Omega\left(\max\left\{\frac{C_{f}^{2}}{\epsilon_{f}^{2}},\frac{C_{g}^{2}}{\epsilon_{g}^{2}}\right\}D^{2}\right)$
@@
Any feasible solution to Problem (1) is also a feasible solution to Problem (4) as the optimal solution ${\bf{x}}^{*}$ satisfies ${\bf{x}}^{*}\in\mathcal{B}({\bf{x}}_{0},D)$, thus $\hat{f}^{*}\leq f^{*}$. For any ${\bf{x}}$ that satisfies
@@
$\psi^{*}(t+\Delta)$ $=\min_{{\bf{x}}\in{\mathcal{B}}({\bf{x}}_{0},D)}\{\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})+\Delta\}\}-\Delta$
$\geq\min_{{\bf{x}}\in{\mathcal{B}}({\bf{x}}_{0},D)}\{\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}\}-\Delta$
@@
Suppose that $t<\hat{f}^{*}$ and $\psi^{*}(t)\leq 0$, then there exists a $\hat{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)$ such that $\tilde{g}(\hat{\bf{x}})\leq 0,f(\hat{\bf{x}})-t\leq 0$. Then $\hat{\bf{x}}$ is a feasible solution to Problem (4), with $f(\hat{\bf{x}})\leq t<\hat{f}^{*}$, contradiction to the fact that $\hat{f}^{*}$ is the optimal value to Problem (4). Thus for any $t<\hat{f}^{*}$, it holds that $\psi^{*}(t)>0$.
@@
Theorem 3.2 in Bubeck et al. [4] states that applying subgradient method to any $C$-Lipschitz convex function $h$ on a compact set $Q$ with diameter $R$ gives
$h(\frac{1}{K}\sum_{i=0}^{K-1}{\bf{x}}_{i})-h^{*}\leq\frac{RC}{\sqrt{K}}.$
Here $h({\bf{x}})=\psi(t,{\bf{x}})=\max(f({\bf{x}})-t,\tilde{g}({\bf{x}}))$,$R=2D,C=\max(C_{f},C_{g})$. Then when $K\geq\frac{16D^{2}C^{2}}{\epsilon^{2}}$, it holds that
$\hat{\psi}^{*}(t)=\psi(t,\frac{1}{K}\sum_{i=0}^{K-1}{\bf{x}}_{i})\leq\psi^{*}(t)+\frac{\epsilon}{2}.$
@@
Theorem 2.3.5 in Nesterov [19] originally states that applying genearlized-gradient method to any $\mu$-strong convex and $L$-smooth convex function $h$ gives
@@
$\psi(t,{\bf{x}}_{K})-\psi^{*}\leq\frac{4L}{\frac{1}{2}L(K+1)^{1}}(\frac{L}{2}\|{\bf{x}}_{0}-{\bf{x}}^{*}\|_{2}^{2}+\frac{L}{4}\|{\bf{x}}_{0}-{\bf{x}}^{*}\|_{2}^{2})\leq\frac{24LD^{2}}{K^{2}}.$
Then when $K\geq\sqrt{\frac{48L}{\epsilon}}D$, it holds that
@@
${\bf{x}}_{1}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}\bar{f}(t,{\bf{x}};{\bf{y}}_{k}),$
${\bf{x}}_{2}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla g({\bf{y}}_{k})\right\|=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)}\bar{g}(t,{\bf{x}};{\bf{y}}_{k}).$
@@
Note that $\mathcal{H}$ is a hyperplane. Since ${\bf{x}}_{k+1}$ minimizes $\bar{\psi}(t,{\bf{x}};{\bf{y}}_{k})$ over $\mathcal{B}({\bf{x}}_{0},D)$,it also minimizes $\bar{\psi}(t,{\bf{x}};{\bf{y}}_{k})$ over the convex subset $\mathcal{B}({\bf{x}}_{0},D)\cap\mathcal{H}$. Given that $\bar{f}(t,{\bf{x}};{\bf{y}}_{k})=\bar{g}({\bf{x}};{\bf{y}}_{k})$ in this subset, we have
${\bf{x}}_{k+1}$ $=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)\cap\mathcal{H}}\bar{\psi}(t,{\bf{x}};{\bf{y}}_{k})$
$=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)\cap\mathcal{H}}\bar{f}(t,{\bf{x}};{\bf{y}}_{k})$
$=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{B}({\bf{x}}_{0},D)\cap\mathcal{H}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|$
@@
Both FC-BiOLip and FC-BiOsm satisfy Assumption 3.4.
@@ E.1 Problem (10)
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\;\frac{1}{2}\|{\bf{x}}\|_{2}^{2}\quad\text{s.t.}\quad A{\bf{x}}={\bf{b}}.$
@@
To be fair, all algorithms start from a random point ${\bf{x}}_{0}$ of unit length as the initial point. For our Algorithm 1, we take a slightly different implementation, that instead of setting the maximum number of iterations of the inner subroutine to be $T^{\prime}=T/N$, we preset $T^{\prime}=8000$. If current ${\bf{x}}_{k}$ already satisfies $\psi(t,{\bf{x}}_{k})\leq\frac{\epsilon}{2}$, then terminate the inner subroutine directly. We adopt the warm-start strategy as described in Appendix B.2. We set $D=2$ and we set $L=0$ since $f({\bf{x}})$ is nonnegative. For FC-BiOLip, we set $\eta=3\times 10^{-4}$. For AGM-BiO, we set $\gamma=1$. For PB-APG, we set $\gamma=10^{5}$. For Bi-SG, we set $\alpha=0.75$ and $c=\frac{1}{L_{f}}$. For a-IRG, we set $\eta_{0}=10^{-3}$ and $\gamma_{0}=10^{-3}$. For CG-BiO, we set $\mathcal{Z}=\mathcal{B}({\bf{x}}_{0},D)$ and $\gamma_{k}=0.5/(k+1)$. For Bisec-BiO, we set the maximum number of iterations of the internal APG process to be $T^{\prime}=10000$. For FC-BiOsm, FC-BiOLip and Bisec-BiO, solving $\hat{g}^{*}$ takes $15000$ iterations; for CG-BiO, solving $\hat{g}^{*}$ takes $10000$ iterations. The results of such pretreatments are also plotted in Figure 1.
We remark that the output of our method satisfies that $f(\hat{\bf{x}})<f^{*}$. Thus although $|f(\hat{\bf{x}})-f^{*}|>\epsilon_{f}$, indeed a $(\epsilon_{f},\epsilon_{g})$-weak optimal solution is solved by FC-BiOsm.
@@ E.2 Problem (11)
In this experiment, a sample of $5000$ instances is taken from the “rcv1.binary” dataset [15, 8] as the training set $(A^{tr},{\bf{b}}^{tr})$; another $5000$ instances is sampled as the validation set $(A^{val},{\bf{b}}^{val})$. Each label ${\bf{b}}^{tr}_{i}$ (or ${\bf{b}}^{val}_{i}$) is either $+1$ or $-1$. To the best of our knowledge, no existing solver could obtain the exact value of $f^{*}$ and $g^{*}$ of Problem (11). Thus we only plot function value $f({\bf{x}}_{k})$ and $g({\bf{x}}_{k})$, instead of suboptimality.
@@
In this experiment, we set the initial point ${\bf{x}}_{0}=0$ for all methods. The implementation of our Algorithm 1 is similar to that in the first experiment. We set the maximum number of iterations of the subroutine to be $T^{\prime}=500$ and $T^{\prime}=1000$ for FC-BiOsm and FC-BiOLip respectively. We set $D=300$. For FC-BiOLip, we set $\eta=2$. For AGM-BiO, we set $\gamma=1/(\frac{2L_{g}}{L_{f}}T^{\frac{2}{3}}+2)$ as in [6, Theorem 4.4]. For PB-APG, we set $\gamma=10^{4}$. For Bi-SG, we set $\alpha=0.75$ and $c=\frac{1}{L_{f}}$. For a-IRG, we set $\eta_{0}=10^{3}$ and $\gamma_{0}=0.1$. For CG-BiO, we set $\mathcal{Z}=\mathcal{B}({\bf{x}}_{0},D)$ and we set $\gamma_{k}=\frac{2}{k+2}$. For FC-BiOsm, solving $\hat{g}^{*}$ takes $1000$ iterations; for FC-BiOLip and CG-BiO, solving $\hat{g}^{*}$ takes $1500$ iterations. As in the first experiment, the results of such pretreatments are also plotted in Figure 2.

### v2

@@ Appendix C Proofs for Section 5
which implies that any zero-respecting first-order method needs at least
$T=\Omega\left(\frac{C_{f}^{2}}{\epsilon_{f}^{2}}D^{2}\right)=\Omega\left(\max\left\{\frac{C_{f}^{2}}{\epsilon_{f}^{2}},\frac{C_{g}^{2}}{\epsilon_{g}^{2}}\right\}D^{2}\right)$
If $\frac{C_{f}}{\epsilon_{f}}\leq\frac{C_{g}}{\epsilon_{g}}$, we set the upper-level and lower-level objective $f,g:{\mathbb{R}}^{T}\to{\mathbb{R}}$ be:
$f({\bf{x}})=0,\quad g({\bf{x}})=r_{T,C_{g},D}({\bf{x}}).$
By similar arguments, any zero-respecting first-order method needs at least
$T=\Omega\left(\frac{C_{g}^{2}}{\epsilon_{g}^{2}}D^{2}\right)=\Omega\left(\max\left\{\frac{C_{f}^{2}}{\epsilon_{f}^{2}},\frac{C_{g}^{2}}{\epsilon_{g}^{2}}\right\}D^{2}\right)$
@@
Any feasible solution to Problem (1) is also a feasible solution to Problem (4), thus $\hat{f}^{*}\leq f^{*}$. For any ${\bf{x}}$ that satisfies
@@
$\psi^{*}(t+\Delta)$ $=\min_{{\bf{x}}\in{\mathcal{Z}}}\{\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})+\Delta\}\}-\Delta$
$\geq\min_{{\bf{x}}\in{\mathcal{Z}}}\{\max\{f({\bf{x}})-t,\tilde{g}({\bf{x}})\}\}-\Delta$
@@
Suppose that $t<\hat{f}^{*}$ and $\psi^{*}(t)\leq 0$, then there exists a $\hat{\bf{x}}\in\mathcal{Z}$ such that $\tilde{g}(\hat{\bf{x}})\leq 0,f(\hat{\bf{x}})-t\leq 0$. Then $\hat{\bf{x}}$ is a feasible solution to Problem (4), with $f(\hat{\bf{x}})\leq t<\hat{f}^{*}$, contradiction to the fact that $\hat{f}^{*}$ is the optimal value to Problem (4). Thus for any $t<\hat{f}^{*}$, it holds that $\psi^{*}(t)>0$.
@@
Theorem 3.2 in Bubeck et al. [4] states that applying subgradient method to any $C$-Lipschitz convex function $h$ on a compact set $Q$ with diameter $D$ gives
$h\left(\frac{1}{K}\sum_{i=0}^{K-1}{\bf{x}}_{i}\right)-h^{*}\leq\frac{DC}{\sqrt{K}}.$
Here $h({\bf{x}})=\psi(t,{\bf{x}})=\max(f({\bf{x}})-t,\tilde{g}({\bf{x}}))$, $C=\max(C_{f},C_{g})$. Then when $K\geq\frac{4D^{2}C^{2}}{\epsilon^{2}}$, it holds that
$\hat{\psi}^{*}(t)=\psi\left(t,\frac{1}{K}\sum_{i=0}^{K-1}{\bf{x}}_{i}\right)\leq\psi^{*}(t)+\frac{\epsilon}{2}.$
@@
Theorem 2.3.5 in Nesterov [19] originally states that applying genearlized-gradient method to any $\mu$-strongly convex and $L$-smooth convex function $h$ gives
@@
$\psi(t,{\bf{x}}_{K})-\psi^{*}\leq\frac{4L}{\frac{1}{2}L(K+1)^{2}}\left(\frac{L}{2}\|{\bf{x}}_{0}-{\bf{x}}^{*}\|_{2}^{2}+\frac{L}{4}\|{\bf{x}}_{0}-{\bf{x}}^{*}\|_{2}^{2}\right)\leq\frac{6LD^{2}}{K^{2}}.$
Then when $K\geq\sqrt{\frac{12L}{\epsilon}}D$, it holds that
@@
${\bf{x}}_{1}=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\bar{f}(t,{\bf{x}};{\bf{y}}_{k}),$
${\bf{x}}_{2}=\operatorname*{arg\,min}_{{\bf{x}}\in{\mathcal{Z}}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla g({\bf{y}}_{k})\right\|=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}}\bar{g}(t,{\bf{x}};{\bf{y}}_{k}).$
@@
Given that $\bar{f}(t,{\bf{x}};{\bf{y}}_{k})=\bar{g}({\bf{x}};{\bf{y}}_{k})$ in the subset $\mathcal{Z}\cap\mathcal{H}$, we have
${\bf{x}}_{k+1}$ $=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}\cap\mathcal{H}}\bar{\psi}(t,{\bf{x}};{\bf{y}}_{k})$
$=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}\cap\mathcal{H}}\bar{f}(t,{\bf{x}};{\bf{y}}_{k})$
$=\operatorname*{arg\,min}_{{\bf{x}}\in\mathcal{Z}\cap\mathcal{H}}\left\|{\bf{x}}-{\bf{y}}_{k}+\frac{1}{L}\nabla f({\bf{y}}_{k})\right\|$
@@
Both FC-BiOLip and FC-BiOsm on the domain ${\mathcal{B}}(\mathbf{{\bf{x}}}_{0},D)$ satisfy Assumption 3.4.
@@
Remark C.1.
By similar analysis, we can show that the conditional gradient type methods for simple bilevel problems [12, 5] on domain ${\mathcal{B}}({\bf{x}}_{0},D)$ also fall into the zero-respecting function class. At each iteration $k$, their methods relies on the following linear program as an oracle.
${\bf{s}}_{k}=\arg\min_{{\bf{s}}\in{\mathcal{X}}_{k}}\langle\nabla f({\bf{x}}_{k},{\bf{s}}\rangle,$
where ${\mathcal{X}}_{k}=\left\{{\bf{s}}\in{\mathcal{Z}}:\langle\nabla g({\bf{x}}_{k}),{\bf{s}}-{\bf{x}}_{k}\rangle\leq g({\bf{x}}_{0})-g({\bf{x}}_{k})\right\}$. According to our proof of Proposition C.1, it suffices to prove that ${\bf{s}}_{k}$ is a linear combination of $\nabla f({\bf{x}}_{K})$ and $\nabla g({\bf{x}}_{k})$ then the remaining proofs are the same. Similarly, this can be seen by the KKT condition when ${\mathcal{Z}}={\mathcal{B}}(\mathbf{0},D)$, which is
$\nabla f({\bf{x}}_{k})+2\lambda{\bf{x}}+\mu\nabla g({\bf{x}}_{k})=0,$
where $\lambda\geq 0$ and $\mu\geq 0$ are dual variables.
@@ E.1 Problem (11)
$\min_{\mathbf{x}\in{\mathcal{Z}}}\;\frac{1}{2}\|{\bf{x}}\|_{2}^{2}\quad\text{s.t.}\quad A{\bf{x}}={\bf{b}}.$
@@
To be fair, all algorithms start from a random point ${\bf{x}}_{0}$ of unit length as the initial point. For our Algorithm 1, we take a slightly different implementation, that instead of setting the maximum number of iterations of the inner subroutine to be $T^{\prime}=T/N$, we preset $T^{\prime}=8000$. If current ${\bf{x}}_{k}$ already satisfies $\psi(t,{\bf{x}}_{k})\leq\frac{\epsilon}{2}$, then terminate the inner subroutine directly. We adopt the warm-start strategy as described in Appendix B.2. We set $L=0$ since $f({\bf{x}})$ is nonnegative. For FC-BiOLip, we set $\eta=3\times 10^{-4}$. For AGM-BiO, we set $\gamma=1$ as in [6, Theorem 4.1]. For PB-APG, we set $\gamma=10^{5}$. For Bi-SG, we set $\alpha=0.75$ and $c=\frac{1}{L_{f}}$. For a-IRG, we set $\eta_{0}=10^{-3}$ and $\gamma_{0}=10^{-3}$. For CG-BiO, we set $\gamma_{k}=0.5/(k+1)$. For Bisec-BiO, we set the maximum number of iterations of the internal APG process to be $T^{\prime}=10000$. For FC-BiOsm, FC-BiOLip and Bisec-BiO, solving $\hat{g}^{*}$ takes $15000$ iterations; for CG-BiO, solving $\hat{g}^{*}$ takes $10000$ iterations. The results of such pretreatments are also plotted in Figure 1.
@@ E.2 Problem (12)
In this experiment, a sample of $5000$ instances is taken from the “rcv1.binary” dataset [15, 8] as the training set $(A^{tr},{\bf{b}}^{tr})$; another $5000$ instances is sampled as the validation set $(A^{val},{\bf{b}}^{val})$. Each label ${\bf{b}}^{tr}_{i}$ (or ${\bf{b}}^{val}_{i}$) is either $+1$ or $-1$.
@@
In this experiment, we set the initial point ${\bf{x}}_{0}=0$ for all methods. The implementation of our Algorithm 1 is similar to that in the first experiment. We set the maximum number of iterations of the subroutine to be $T^{\prime}=500$ and $T^{\prime}=1000$ for FC-BiOsm and FC-BiOLip respectively. For FC-BiOLip, we set $\eta=2$. For AGM-BiO, we set $\gamma=1/(\frac{2L_{g}}{L_{f}}T^{\frac{2}{3}}+2)$ as in [6, Theorem 4.4]. For PB-APG, we set $\gamma=10^{4}$. For Bi-SG, we set $\alpha=0.75$ and $c=\frac{1}{L_{f}}$. For a-IRG, we set $\eta_{0}=10^{3}$ and $\gamma_{0}=0.1$. For CG-BiO, we set $\gamma_{k}=\frac{2}{k+2}$. For FC-BiOsm, solving $\hat{g}^{*}$ takes $1000$ iterations; for FC-BiOLip and CG-BiO, solving $\hat{g}^{*}$ takes $1500$ iterations. As in the first experiment, the results of such pretreatments are also plotted in Figure 2.